### 19 Apr 2026 v1

# Classical Percolation from Quantum Metric in Flat-Band Delocalization

Abstract
The quantum metric is a fundamental ingredient of band quantum geometry and has recently attracted intense interest, with most of its transport signatures appearing in the intrinsic second order nonlinear conductivity. In the clean limit, previous works argued that linear response conductivity is insensitive to the quantum metric, while the Berry curvature yields an intrinsic anomalous Hall contribution. Here we combine analytic derivations with new numerics to show that disorder modifies the linear response conductivity dominated by geometric conductivity which is determined by the real space quantum metric. Focusing on a two dimensional multi-flatband stub-pyrochlore lattice, we identify a critical delocalized regime sandwiched between flat band localization and Anderson localization, characterized by finite geometric conductivity. Upon including spin orbit coupling, this regime evolves into a diffusive metallic phase, constituting a two dimensional inverse Anderson transition. Moreover, exploiting the connection between the real space quantum metric marker and the Wannier function spread, we construct a bond-percolation model on a square lattice. The resulting percolation region quantitatively coincides the critical delocalized regime, the exponent of which supports a classical percolation universality class. These findings suggest that flat band delocalization can be understood as a classical percolation of quantum metric puddles. This advances our understanding of quantum geometric contributions to transport and establishes linear response measurements as a new avenue for accessing the quantum metric.
The quantum metric provides a gauge invariant measure of the “distance” between neighboring Bloch states in Hilbert space. It constitutes the real part of the quantum geometric tensorProvost and Vallée (1980); Resta (2011) of which the imaginary part is the Berry curvatureBerry (1984); Simon (1983); Aharonov and Anandan (1987). Beyond its formal definition, the quantum metric has become a unifying ingredient across disparate phenomenaTörmä et al. (2022); Yu et al. (2025); Liu et al. (2024): it contributions to the superfluid weight in flat band superconductorsPeotta and Törmä (2015); Julku et al. (2016); Liang et al. (2017); Hofmann et al. (2020); Herzog-Arbeitman et al. (2022), in strongly correlated flat band settings it supplies geometric structure that constrains and stabilizes fractional Chern insulatorRoy (2014); Mera and Ozawa (2021); Ledwith et al. (2023).
In transport, a key advance is that the quantum metric can drive intrinsic nonlinear conductivitiesKaplan et al. (2024); Wang et al. (2021); Liu et al. (2021) that are not controlled by a relaxation time, making it an experimentally attractive quantum geometric channel complementary to Berry curvature mechanisms. Such intrinsic quantum metric signatures are particularly transparent in antiferromagnetsGao et al. (2023); Wang et al. (2023) that break inversion P and time-reversal T individually while preserving the combined PT symmetry which enforces vanishing Berry curvatureSodemann and Fu (2015). However, previous theoretical studiesHuhtinen and Törmä (2023); Verma and Queiroz (2025) have shown that in the clean limit, the linear conductivity receives no contribution from the quantum metric, with only the Berry curvature contributing to the anomalous Hall conductanceXiao et al. (2010).
The quantum metric can become the leading geometric scale controlling measurable response coefficientsChen and Law (2024) in flat band systems, where conventional dispersion driven channels are strongly suppressed. There are two types of flat bands generated by destructive interferenceRhim and Yang (2019). The first type, observed in systems such as 2D Lieb, 2D pyrochlore, and Kagome lattices, forces the flat band to touch the edge of a dispersive band, resulting in “protected” crossingsBergman et al. (2008) due to the real-space topology of the Compact Localized States (CLS)Sutherland (1986); Maimaiti et al. (2017); Röntgen et al. (2018); Aoki et al. (1996). The second type features flat bands whose position relative to dispersive bands is tunable, as in stub and cross-stitch lattices. The flat band in the first case shares similar localization properties with the dispersive bands, whereas in the second case, it differs, indicating that novel transport in flat bands can only be realized in such systemsLeykam et al. (2017).
Many works have investigated the localization properties in flat band systems induced by destructive interferenceLeykam et al. (2013); Bodyfelt et al. (2014); Vicencio and Johansson (2013); Ahmed et al. (2022); Hou et al. (2025), especially the theoretical and experimental discovery of inverse Anderson transitionsGoda et al. (2006); Li et al. (2022); Zhang et al. (2023). Recent theoretical studies have linked diffusion transport in one dimensional flat band and twsit bilayer graphene to quantum metricGuerrero et al. (2025); Chau et al. (2024). Notably, both theoretical predictions and experiments on superconducting qubit arrays reveal disorder induced delocalization in multi-flat band systems, with a transition from flat band localization to Anderson localizationAnderson (1958); Evers and Mirlin (2008); Qi et al. (2025), whereas dispersive bands do not show this phenomenonRosen et al. (2025). This observation highlights the necessity of studying novel transport properties in multi-flat band systems affected by destructive interference, in order to understand the influence of quantum metric.
In this Letter, we demonstrate that flat band delocalization with linear conductivity is attributed to a classical percolation process governed by the quantum metric. We present a detailed analytic derivation of a linear geometric conductance formula that captures the quantum metric contribution in disordered flat bands. It establishes a direct connection between transport and real space quantum metricBianco and Resta (2011). We then adopt a two dimensional multi-flatband stub-pyrochlore model designed to exclude mechanisms other than the quantum metric, and introduce disorder in the energy window between two flat bands. With spin orbit coupling (SOC) and without SOC, both the purely geometric conductance and the conductance computed from transport simulations are obtained. Finally, by using real space quantum metric marker as a component, we construct a classical geometric percolation model that accounts for the emergence of the geometric conductance.
The real space quantum metric de Sousa et al. (2023); Marrazzo and Resta (2019); Marsal and Black-Schaffer (2024); Roche Carrasco et al. (2025); Avdoshkin et al. (2025); Romeral et al. (2025) restricted to the occupied states is:
[…]
where $\hat{P}$ and $\hat{Q}$ are occupied and unoccupied projectors, $\hat{r}_{\mu}$ is the position operator, all in the band basis.
We define the energy difference Lorentzian distribution in the band basis as
$L_{mn}(\omega)\;\equiv\;\frac{2\eta\omega_{mn}}{\omega_{mn}^{2}+\eta^{2}},\qquad(m\neq n),$ (2)
with $L_{nn}(\omega)=0$, where $\eta$ is the band broadening, vanishes in the clean limit, $\omega_{mn}$ is the energy difference between band $m$ and $n$.
The geometric conductance is,
$\sigma_{\mu\nu}^{\rm geo}$ $=\frac{e^{2}}{Vh}\mathrm{Tr}\!\left(g_{\mu\nu}\,L\right),$ (3)
detailed derivations are provided in the supplementary material.
Figure 1: 2D stub-pyrochlore lattice. (a) Schematic of the two dimensional multi-flatband stub–pyrochlore model on a square lattice, illustrating the hopping processes $t_{x},t_{y}$, the crossing couplings $t_{c1},t_{c2}$, and the stub hybridizations $\lambda_{x},\lambda_{y}$. (b) Band structure of the SOC-free Hamiltonian along the high symmetry path $X\!-\!\Gamma\!-\!Y$. (c) Band structure of the corresponding spinful Hamiltonian with Rashba SOC $H_{R}$ along the same path.
Previous studiesBouzerar and Mayou (2021); Mitscherling and Holder (2022); Mera and Mitscherling (2022) have suggested that a finite $\eta$, potentially arising from disorder or finite temperature effects, can induce nontrivial transport features in flat bands. However, these works remain largely qualitative and lack quantitative calculations of $\eta$ with concrete numerical values. Here, using an effective medium formulation within the self consistent Born approximationGroth et al. (2009), we attribute the disorder contribution to a self consistently determined self energy,
$\Sigma=\frac{W^{2}}{12}\frac{a^{2}}{4\pi^{2}}\int_{BZ}d^{2}\mathbf{k}\frac{1}{E_{f}+i\eta-H(\mathbf{k})-\Sigma}$ (4)
where $W$ is the disorder strength, a randomon-site disorder potential is uniformly distributed in the range $\frac{W}{2},\frac{W}{2}]$. The imaginary part of the self energy components directly yields the disorder induced spectral broadeningVu and Das Sarma (2022); Shen et al. (2024), $\eta\sim-\frac{1}{N_{b}}\,\operatorname{Im}tr\Sigma^{R}$, $N_{b}$ is the number of bands.
To obtain flat band transport behavior distinct from that of dispersive bands and to extract an unambiguous conductance contribution driven purely by the quantum metric, we employ a model with two CLSs in a 2D stub-pyrochlore latticeNeves et al. (2024) in Fig.1(a), where the square has side length 0.5. The Hamiltonian is defined by:
$H(\mathbf{k})$ $=\begin{pmatrix}\epsilon_{A_{1}}&-t_{x}f_{x}&-\lambda_{1}&t_{c1}&-t_{y}f_{y}&0\\ -t_{x}f_{x}&\epsilon_{B_{1}}&0&-t_{y}f_{y}&t_{c2}&0\\ -\lambda_{1}&0&\epsilon_{S_{1}}&0&0&0\\ t_{c1}&-t_{y}f_{y}&0&\epsilon_{A_{2}}&-t_{x}f_{x}&-\lambda_{2}\\ -t_{y}f_{y}&t_{c2}&0&-t_{x}f_{x}&\epsilon_{B_{2}}&0\\ 0&0&0&-\lambda_{2}&0&\epsilon_{S_{2}}\end{pmatrix}$. (5)
where $f_{x}=2\cos\!\left(\frac{k_{x}}{2}\right),f_{y}=2\cos\!\left(\frac{k_{y}}{2}\right),t_{x}=1.0,t_{y}=1.0,$ $\epsilon_{S_{1}}$ = 0.546, $\epsilon_{S_{2}}$ = 0.115, $\epsilon_{A_{1}}$ = 0.808, $\epsilon_{A_{2}}$ = 0.927, $\lambda_{1}$ = 0.102, $\lambda_{2}$ = 0.191, $\epsilon_{B_{1}}$ = -0.262, $\epsilon_{B_{2}}$ = 0.014, $t_{c1}$ = 0.239, $t_{c2}$ = 0.100. The CLSs hybridize and open a trivial gap. Further details on the Hamiltonian construction and parameter generation are provided in the supplementary material. In Fig. 1(b), we present the high symmetry line band structure of the model, where two flat bands, arising from destructive interference, are observed at $E_{f}=0.1$ and $E_{f}=0.54$, well separated from the dispersive bands among the six energy bands. In Fig. 1(c), we incorporate the spin component and adding a Rashba SOCRashba (1960); Bychkov and Rashba (1984) term,
$H_{\mathrm{R}}(k)=-2\alpha_{\mathrm{R}}\big(\sin k_{y}\,\sigma_{x}-\sin k_{x}\,\sigma_{y}\big)$ (6)
We choose $\alpha_{\mathrm{R}}=0.5$. The flat bands at $E_{f}=0.1$ and $E_{f}=0.54$ remain unaffected.
Then we investigate the transport between the two flat bands at $E_{f}=0.3$. The system consists of an $L\times L$ sample connected to two conducting leads. To reduce contact resistance, the leads are chosen to have the same Hamiltonian as the sample, but with the Fermi energy $E_{f}=-0.5$, placing the leads in the conducting regime. We examine the effects of onsite disorder on the square lattice, where no disorder on the stub. The disorder is modeled by randomly choosing values from a uniform distribution over the range [W/2,W/2][-W/2,W/2], with $W$ being the disorder strength and is averaged over 2000 samples. The two-terminal conductance $G$ is computed using the Landauer–Buttiker formulaLandauer (1970); Büttiker (1988); Fisher and Lee (1981); Datta (1995) and the recursive Green’s function methodMacKinnon (1985); Sancho et al. (1985).
$G=\frac{e^{2}}{h}\operatorname{Tr}\left[\Gamma_{\mathrm{l}}\mathrm{G}^{\mathrm{r}}\Gamma_{\mathrm{r}}\mathrm{G}^{\mathrm{a}}\right],$ (7)
where $\Gamma_{\mathrm{l}},\Gamma_{\mathrm{r}}$ are the line width function describing the coupling to the left (l) or right (r) lead. $\mathrm{G}^{\mathrm{r}},\mathrm{G}^{\mathrm{a}}$ are the retarded and advanced Green’s function of the scattering region. The trace accounts for contributions from all transport channels.
Figure 2: Two terminal transport calculations and geometric conductance simulations. (a) Disorder induced spectral broadening $\eta$ as a function of disorder strength $W$ for both the SOC-free model and the Rashba-SOC model; the dashed horizontal line indicates the band gap $\Delta$. (b) SOC-free case: two-terminal conductance $G$ and geometric conductance $\sigma_{\rm geo}$ versus $W$ for different system sizes $L$; dashed curves denote the total conductance obtained from Green’s function calculation, while symbols show $\sigma_{\rm geo}$ obtained from eq.3. (c) Rashba-SOC case: $G$ and $\sigma_{\rm geo}$ versus $W$ for different $L$, an intermediate diffusive-metal regime emerges. (d) Rashba-SOC case: 2000 sample averaged $\langle\log G\rangle$ versus $W$ for different $L$; inset: finite size scaling near the first crossing yields the critical exponent $\nu$.
The scattering rate $\eta$ for the model without SOC and with Rashba SOC are shown in Fig.2 (a). In the weak disorder regime near the clean limit, $\eta$ remains zero. Upon increasing it turns on beyond a threshold and then grows monotonically with disorder strength. For $W\sim 2$, the disorder-induced broadening becomes comparable to the gap, indicating that the system has entered the regime where quantum metric dominatesBouzerar and Mayou (2021); Mitscherling and Holder (2022); Mera and Mitscherling (2022).
In Fig. 2(b), both transport conductance and geometric conductivity of model without SOC are employed, and we show the conductance as a function of disorder strength. In the clean limit, the conductance is almost zero, corresponding to the insulating state caused by flat band localization. Notably, for $W>2$, the conductance increases rapidly to the order of $e^{2}/h$. The peak occurs at $W\sim 3.4$, after which the conductance starts to decrease, eventually entering the insulating phase of Anderson localization. The points shown in the figure represent the geometric conductivity, calculated using eq.3, which closely matches the transport conductance. As the system size increases, the conductance decreases slowly and monotonically. This behavior aligns with the predictions of single parameter scaling theoryAbrahams et al. (1979). Previous work on planar pyrochlore lattices with weak disorder has found critical behavior accompanied by multifractal featuresChalker et al. (2010); Evers and Mirlin (2008), where the system is neither Anderson localized nor spatially extended. Notably, a recent study demonstrated a strong enhancement of the space resolved quantum metric in a disordered setting and traced its origin to multifractal critical wave functionsMarsal and Black-Schaffer (2024), consistent with our findings. These demonstrate that the conductance observed under disorder is primarily due to the quantum metric, with no corresponding behavior in the clean limit.
In Fig. 2(c), we present the model that includes Rashba SOC. The resulting conductance versus disorder strength curves show a similar trend to that in Fig. 2(b), with flat band localization occurring at $W\leq 2.2$ and Anderson localization at $W\geq 6.5$. However, the conductance values are larger, particularly in the intermediate region where two critical points emerge. In the range $2.2\leq W\leq 6.5$, as the system size increases, the conductance grows, exhibiting metallic behavior. The geometric conductivity calculated using eq.3 closely matches the transport conductance, and the observed larger conductance and metallic behavior can be attributed to the strengthening of the quantum metric due to Rashba SOCSala et al. (2025). To investigate the metallic phase and extract the critical exponent, we calculate the sample averaged log(G)Slevin et al. (2001) and plot it as a function of disorder strength W for various system sizes in Fig. 2(d). The curves reveal two crossing points, signaling a two dimensional inverse Anderson transitionGoda et al. (2006). We fit the critical exponent for the transition from flat band localization to the diffusive metal phase to be approximately 2.24, close to previous studies of 2D symplectic metal-insulator transitionsHikami et al. (1980); Asada et al. (2002, 2004).
To elucidate the microscopic origin of the geometric conductivity observed in the critical window in the absence of SOC, we employ insulating leads by setting the lead Fermi energy to $E_{f}$=0.3, matching the central region, and evaluate the disorder dependence of the sample averaged log(G). In Fig. 3(a), we find a pronounced intermediate regime between flat band localization and Anderson localization where averaged log(G) grows slowly with increasing system size, signaling quasi-metallic behaviorAbrahams et al. (2001). Importantly, because the leads are insulating, this response cannot be attributed to carriers injected from the leads; rather, it reflects spontaneously transport sustained by the intrinsic electronic states of the disordered flat band system, naturally suggesting a percolation type delocalization scenario. Fitting the two critical points $W_{c1}$ and $W_{c2}$ yields $\nu_{1}\sim$1.38 and $\nu_{2}\sim$1.32, close to the two dimensional classical percolation exponent $\nu\sim$1.3Shashkin et al. (1994); Dunford et al. (2001).
Figure 3: Geometric percolation model. (a) SOC-free model with insulating leads: 3000 sample averaged $\langle\log G\rangle$ as a function of disorder strength $W$ for different system sizes $L$. Inset: finite-size scaling analysis in the vicinity of the two crossing points, yielding the critical exponents $\nu_{1}$ and $\nu_{2}$. (b) Schematic of the geometric percolation picture. Colored hemispheres represent effective potential valleys (“puddles”) whose local extent encodes the spread. Neighboring valleys are connected when quantum tunneling is sufficiently strong; in the illustration, the bright links highlight a system-spanning conducting path formed by connected valleys.
To further substantiate this picture, we introduce a geometric percolation modelKirkpatrick (1973); Isichenko (1992); Stauffer and Aharony (1994) rooted in real space quantum metric marker in Fig. 3(b). The gauge invariant Wannier spread $\Omega_{I}$, which quantifies the variance of the real space charge distributionResta and Sorella (1999), can be expressed in terms of a real space quantum metric marker gMarzari and Vanderbilt (1997); Cárdenas-Castillo et al. (2024),
$\Omega_{I}=\frac{V_{\text{cell }}}{\hbar^{D-2}}\operatorname{Tr}g_{\mu\nu}$
we thus associate each effective potential valley supporting a compact localized orbit with a site of an auxiliary lattice, and connect nearest-neighbor valleys by links that represent quantum tunneling between themMeir (1999, 2000). When disorder is weak, $g_{i}$ is small and electronic trajectories remain trapped inside individual valleys, with only weak tunneling across the saddle region separating adjacent valleys. As disorder increases, the destructive interference confinement is progressively lifted, which enhances the gauge-invariant spread between neighboring valleys. Once this enhancement exceeds the local saddle-point threshold, two previously isolated trajectories coalesce: the electron can move freely from one valley to its neighbor, and the link connecting them becomes a perfect linkDubi et al. (2005b). In this coarse-grained network, joining two sites by a perfect link effectively merges them into a single site. Consequently, only when the perfect links percolate and form a system spanning cluster does the transmission become length independent, so that the critical percolation threshold is well defined and the conductance does not decay with system size. The local saddle-point threshold is model dependent, but the critical exponent is universal. This percolation of perfect links therefore provides a natural geometric mechanism for the localization–delocalization transition observed in transport. The resulting criticality is governed by quantum transport on this emergent network, where tunneling and interference determine the scaling behavior.
For model without SOC, we use the disorder strength $W$, rather than the fraction of perfect links $p$, as the control parameter for the percolation analysis. This choice is motivated by the fact that $p(W)$ is typically analytic in the vicinity of $W_{c}$ and is dominated by its linear term, widely used in the Interger Quantum Hall literatureHuckestein (1995); Kramer et al. (2005). We define the site “spread” as the arithmetic square root of the local marker $g_{i}$ and choose $2\pi$ as the local saddle-point threshold. A nearest-neighbor bond ⟨ij⟩ is a perfect link if their combined spread exceeds the saddle threshold, and is otherwise closed. We generate $N_{s}=200$ disorder realizations for each $(W,L)$, and define the spanning probability $p_{t}$ Kohn (1964); Ziff (1992) as the fraction of samples that admit a conducting path connecting the left and right boundaries. The continuous evolution of $p_{t}$ from 0 to $1$ signals a percolation transition.
Figure 4: Real space quantum metric controlled percolation transition. Spanning probability $p_{t}$ versus disorder strength $W$ for different system sizes $L$, averaged over $N_{s}=200$ disorder realizations. Insets show the same data plotted versus a rescaled horizontal axis ( $z=(W-W_{c})L^{1/\nu}$) to demonstrate scaling collapse near criticality. (a) Window around the first threshold, $W\in[1.7,\,2.4]$. (b) Window around the second threshold, $W\in[4.1,\,5.2]$.
Fig. 4 shows $p_{t}(W)$ for different system sizes $L$. In Fig. 4(a) , around $W\simeq 2$, $p_{t}$ rises sharply from 0 toward $1$ for all $L$, indicating the onset of percolation. This threshold coincides with the transport defined boundary $W_{c1}$ separating the flat-band localized regime from the quasi-metallic window. In the inset of Fig. 4(a), we perform a finite size scaling collapse using the dimensionless scaling variable $z=(W-W_{c})\,L^{1/\nu_{1}}$Dubi et al. (2005a), and find that the data for different $L$ collapse onto a single universal curve, providing strong evidence that the transition is governed by a diverging correlation length with exponent $\nu_{1}$.
In Fig. 4(b), around $W\simeq 4.5$, $p_{t}$ decreases from $1$ back to 0 as $W$ increases, indicating that the system leaves the percolating critical regime and re-enters a localized phase. This second threshold matches the transport defined boundary $W_{c2}$ between the quasi-metallic regime and the Anderson localized regime. The inset of Fig. 4(b) shows an analogous scaling collapse using $z=(W_{c}-W)\,L^{1/\nu_{2}}$ again yielding a single scaling curve. Together, these results support a percolation description of the entire evolution, encoded by the real space quantum metric marker.
In the presence of SOC, the transport can be captured by a quantum tunneling network model in the symplectic symmetry classMinakuchi (1998).
In conclusion, we consider the two dimensional multi-flatband stub-pyrochlore model and identify a pronounced geometric linear conductivity adding disorder between the flat bands, which is directly tied to the real space quantum metric. This region is a critical state, upon incorporating SOC, the enhanced quantum metric transitions from the critical state to a diffusive metallic regime. We model the localization-delocalization transition with a classical percolation framework, where local ”puddles” represent regions of spread, which govern electron transport. The critical state emerges as disorder weakens destructive interference, allowing larger puddles form a percolating cluster, leading to the delocalization transition. As disorder intensifies, puddle sizes shrink, eventually leading to Anderson localization. Therefore, the flat band delocalization can be understood as a classical percolation of quantum metric puddles.
The authors would like to thank Ziyue Qi for the helpful discussions. This work was supported by the National Key R&D Program of China (Grant No. 2023YFA1607400, 2022YFA1403800), the National Natural Science Foundation of China (Grant No.12274436, 11925408, 11921004), the Science Center of the National Natural Science Foundation of China (Grant No. 12188101).

### 21 Apr 2026 v2

@@
-Classical Percolation from Quantum Metric in Flat-Band Delocalization
+Percolation from Quantum Metric in Flat-Band Delocalization
 Abstract
@@
 (2025); Avdoshkin et al.
+(2025); Mitscherling et al.
 (2025); Romeral et al.
@@
 detailed derivations are provided in the supplementary material.
+To exclude contribution from the Drude weightKohn (1964); Scalapino et al.
+(1993); Antebi et al.
+(2024), we computed, in real space under open boundary conditionsBellomia and Resta (2020), the low frequency Drude weight contribution to the conductivity for different disorder strengths, and found that all of them vanish in the dc limit.
+The detailed results are presented in the Supplementary Material.
 Figure 1: 2D stub-pyrochlore lattice.
@@
 (c) Rashba-SOC case: $G$ and $\sigma_{\rm geo}$ versus $W$ for different $L$, an intermediate diffusive-metal regime emerges.
-(d) Rashba-SOC case: 2000 sample averaged $\langle\log G\rangle$ versus $W$ for different $L$; inset: finite size scaling near the first crossing yields the critical exponent $\nu$.
+(d) Rashba-SOC case: 1600 sample averaged renormalized localization length versus $W$ for different $L$; inset: finite size scaling near the first crossing yields the critical exponent $\nu$.
 The scattering rate $\eta$ for the model without SOC and with Rashba SOC are shown in Fig.2 (a).
@@
 (2025).
-To investigate the metallic phase and extract the critical exponent, we calculate the sample averaged log(G)Slevin et al.
-(2001) and plot it as a function of disorder strength W for various system sizes in Fig.
+To investigate the metallic phase and extract the critical exponent, we calculate the sample averaged renormalized localization lengthMacKinnon and Kramer (1981, 1983); Kramer and MacKinnon (1993); Slevin and Ohtsuki (1999) and plot it as a function of disorder strength W for various system sizes in Fig.
 2(d).
@@
 (2006).
-We fit the critical exponent for the transition from flat band localization to the diffusive metal phase to be approximately 2.24, close to previous studies of 2D symplectic metal-insulator transitionsHikami et al.
+We fit the critical exponent for the transition from flat band localization to the diffusive metal phase to be approximately 2.79, close to previous studies of 2D symplectic metal-insulator transitionsHikami et al.
 (1980); Asada et al.
 (2002, 2004).
-To elucidate the microscopic origin of the geometric conductivity observed in the critical window in the absence of SOC, we employ insulating leads by setting the lead Fermi energy to $E_{f}$=0.3, matching the central region, and evaluate the disorder dependence of the sample averaged log(G).
-In Fig.
-3(a), we find a pronounced intermediate regime between flat band localization and Anderson localization where averaged log(G) grows slowly with increasing system size, signaling quasi-metallic behaviorAbrahams et al.
-(2001).
-Importantly, because the leads are insulating, this response cannot be attributed to carriers injected from the leads; rather, it reflects spontaneously transport sustained by the intrinsic electronic states of the disordered flat band system, naturally suggesting a percolation type delocalization scenario.
-Fitting the two critical points $W_{c1}$ and $W_{c2}$ yields $\nu_{1}\sim$1.38 and $\nu_{2}\sim$1.32, close to the two dimensional classical percolation exponent $\nu\sim$1.3Shashkin et al.
-(1994); Dunford et al.
-(2001).
-Figure 3: Geometric percolation model.
-(a) SOC-free model with insulating leads: 3000 sample averaged $\langle\log G\rangle$ as a function of disorder strength $W$ for different system sizes $L$.
-Inset: finite-size scaling analysis in the vicinity of the two crossing points, yielding the critical exponents $\nu_{1}$ and $\nu_{2}$.
-(b) Schematic of the geometric percolation picture.
+Figure 3: Schematic of the geometric percolation picture.
 Colored hemispheres represent effective potential valleys (“puddles”) whose local extent encodes the spread.
@@
 To further substantiate this picture, we introduce a geometric percolation modelKirkpatrick (1973); Isichenko (1992); Stauffer and Aharony (1994) rooted in real space quantum metric marker in Fig.
-3(b).
+3.
 The gauge invariant Wannier spread $\Omega_{I}$, which quantifies the variance of the real space charge distributionResta and Sorella (1999), can be expressed in terms of a real space quantum metric marker gMarzari and Vanderbilt (1997); Cárdenas-Castillo et al.
@@
 4(a), we perform a finite size scaling collapse using the dimensionless scaling variable $z=(W-W_{c})\,L^{1/\nu_{1}}$Dubi et al.
-(2005a), and find that the data for different $L$ collapse onto a single universal curve, providing strong evidence that the transition is governed by a diverging correlation length with exponent $\nu_{1}$.
+(2005a), and find that the data for different $L$ collapse onto a single universal curve, providing strong evidence that the transition is governed by a diverging correlation length with exponent $\nu_{1}$=1.33, which is classicalShashkin et al.
+(1994); Dunford et al.
+(2001).
 In Fig.
@@
 The inset of Fig.
-4(b) shows an analogous scaling collapse using $z=(W_{c}-W)\,L^{1/\nu_{2}}$ again yielding a single scaling curve.
+4(b) shows an analogous scaling collapse using $z=(W_{c}-W)\,L^{1/\nu_{2}}$ with exponent $\nu_{2}$=1.33 again yielding a single scaling curve.
 Together, these results support a percolation description of the entire evolution, encoded by the real space quantum metric marker.
@@
 Therefore, the flat band delocalization can be understood as a classical percolation of quantum metric puddles.
-The authors would like to thank Ziyue Qi for the helpful discussions.
+The authors would like to thank Tobias Holder, Zhida Song, Ziyue Qi for the helpful discussions.
 This work was supported by the National Key R&D Program of China (Grant No.